a +2.9-nC free charge is brought to within 3.1 micrometers of a fixed +12-microC charge and released from rest. if the free charge has a mass of 0.11gms, what is its final velovity when it is very far away?
Answer #1

By Conservation of energy

KQ1Q2/r = (1/2)mV2

(9*109)(2.9*10-9)(12*10-6)/(3.1*10-6)=(1/2)(0.11*10-3)V2

5.5*10-5V2=101.0322

V=1355.34 m/s